OpenStudy (anonymous):

a particle moves along a straight line with equation of motion s=t^6-6t^5. find the value of t at which the acceleration is equal to zero.

OpenStudy (aaronq):

the acceleration is the second derivative of the position equation \(\sf s^{''}(t)=a(t)=30t^4-120t^3=0\) \(\sf t^3(30t-120)=0\) \(\sf t=4\) (not sure what units they are)